# Conformal Prediction Sets Improve Human Decision Making

### v1

@@ 4.2 Experiment Design
Our experimental design was informed by research on the collection of high quality data from crowd sourcing platforms. Mitra et al. (2015) show that the most reliable way to ensure high quality data is by training participants on the task at hand, while providing financial incentives can also be beneficial. Hence, in each experiment participants were asked for consent, given instructions on the task, trained with 20 examples, after which the testing phase began. An example trial screen from ObjectNet is shown in Figure 2, while complete descriptions of the tasks are given in Appendix B. Participants were paid a fixed rate based on the median completion time for their treatment, and were given a financial incentive proportional to their number of correct responses. On average, participants were paid 7.80 GBP/hr.
On each trial the participant was shown a stimulus $x$ and all $M$ class labels, and was forced to classify $x$ as one of the labels. For the top-$k$ and conformal treatments, they were also shown a prediction set from the model, along with the (empirical) coverage guarantee. The correct answer was displayed once the participant confirmed their decision. There was no time limit to enter responses, although for ObjectNet the stimulus was only shown for 0.22s, within the limits of human perception (Fraisse, 1984), to increase the difficulty of the task. The set of $m$ stimuli shown to each participant was drawn from $\mathcal{D}_{\mathrm{test}}$, making each test randomized. Additional details are given in Appendix B.
Figure 2: Main trial screen shown to participants for ObjectNet with conformal set treatment. For ObjectNet only, the stimulus (image) was displayed for 0.22 s. The correct answer is given only after the participant responds.
@@
Figure 4: Human performance (speed) across three tasks and three treatments. Data is shown as mean response taken in seconds to complete all test trials, while error bars show unbiased standard errors ($N=50$).
Table 2: Human Accuracy Comparisons – $p$-values
Comparison ObjectNet GoEmotions Few-NERD
Top-$k$ $>$ Control $p=0.1$ $p=5\mathrm{e-}{5}$ $p=0.003$
Conf. $>$ Control $p=5\mathrm{e-}{5}$ $p=5\mathrm{e-}{9}$ $p=3\mathrm{e-}{8}$
Conf. $>$ Top-$k$ $p=0.01$ $p=0.02$ $p=8\mathrm{e-}{4}$
Table 3: Human Response Time Comparisons – $p$-values
Comparison ObjectNet GoEmotions Few-NERD
Top-$k$ vs. Control $p=0.6$ $p=0.1$ $p=0.3$
Conf. vs. Control $p=0.1$ $p=0.3$ $p=6\mathrm{e-}{5}$
Conf. vs. Top-$k$ $p=0.3$ $p=0.5$ $p=4\mathrm{e-}{5}$
@@ 5.1 Human Performance Measurement
For accuracy, we see that top-$k$ sets mostly do improve human performance compared to the control, while conformal sets lead to better performance than top-$k$ sets. To confirm these visual trends, we conducted pre-planned comparisons between pairs of treatments using $t$-tests as described in Section 3. In all cases except one we reject the null hypothesis that the mean accuracy is the same between treatments, using the significance threshold $p<0.05$ (Table 2, non-significant results in red). Notably, for all three independent tasks we find statistically significant evidence that conformal prediction sets are more useful to humans than top-$k$ sets. Because both types of prediction sets have the same (empirical) coverage, there are only two differences between the methods to which we can ascribe the improvement: conformal sets are smaller on average (Table 1), and conformal sets quantify uncertainty.
We also pre-planned comparisons between the treatments on the average response time. However, in viewing Figure 4 we do not see a consistent trend between all tasks. Conformal sets have the lowest average response time on two tasks, while the control treatment led to the fastest completions on GoEmotions. However, in most cases we do not reject the null hypothesis that the treatment has no effect on response times (Table 3). The additional information provided in the form of prediction sets must be processed by the user and incorporated into decision making, which can outweigh the speed advantages of receiving a curated shortlist.
@@ 5.2 Ablations
Figure 7: Comparison of ObjectNet results between the two models. The ViT-L/14 model achieved 83.3% top-1 accuracy, which was significantly better than ViT-B/32 at 61.1% on ObjectNet test data.
@@
Examining the set size distribution between the two settings helps reveal the source of performance differences. As shown in Figure 5, human performance decreases on examples where prediction sets were larger, but importantly human performance conditioned on set size was comparable between the two settings. By comparison, Figure 6 shows that the large setting generated fewer singleton sets where human accuracy is highest, resulting in the observed overall decrease in human accuracy. Optimizing for set sizes where human performance is highest is one way to improve the usefulness of conformal sets for humans.
@@
By swapping the ViT-L/14 model used in Section 5.1 for a weaker ViT-B/32, we can study the effect of model performance on the usefulness of both top-$k$ and conformal prediction sets. 6(a) shows conformal sets greatly enhanced human performance when using a superior underlying model, whereas for top-$k$ the improved model did not translate to as much human improvement. This is despite the empirical coverage (6(b)) increasing by a similar amount between the treatments (by design of our calibration). Furthermore, the rate at which humans chose an answer from the prediction set, which we call the adoption rate (6(c)), was consistently close to the empirical coverage (coverage was communicated to participants during the test). The outsized increase in accuracy for conformal when the better model was used is a function of the higher quality of sets that it produces; although coverage and adoption rates are the same, humans must still choose the correct answer out of the set when they adopt from it, which is much easier to do for singleton sets. The stronger model produces a distribution of conformal set sizes more skewed towards singleton sets (6(d)), whereas top-$k$ sets always use $k=3$ even with a more confident and correct model.
@@ 5.3 Insights
In Figure 3 we concluded that the combination of human and model (conformal treatment) outperforms humans alone (control) in terms of accuracy. Another relevant question is whether human-model teams outperform the model alone. From Table 1 we see that all treatments outperformed the top-1 accuracy of the model on ObjectNet, but the reverse is true on GoEmotions and Few-NERD. In Figure 10 we examine this comparison in more detail using per-class accuracies for ObjectNet. We see high variance in the model’s per-class accuracy, possibly due to the zero-shot nature of predictions we implemented, whereas humans were more consistent. Notably, low model accuracy (Bandage, Battery, Blanket, Bucket, Cellphone) tended to drag down the human-model team compared to humans alone. This is a point of caution for implementing human-in-the-loop systems, as poor model performance or biases may not be completely overcome by humans. However, when the model outperforms the humans, the human-model team also tends to benefit, sometimes surpassing both individual partners (Backpack, Book, Bottle, Opener, Candle, Sandal). Hence there is a type of ensembling effect (Dietterich, 2000) where one partner’s weaknesses can be corrected by the other.
@@
Participants in the top-$k$ and conformal treatment groups were informed that the prediction sets they were shown contain the true answer with probability $1-\hat{\alpha}$, but were not constrained in how they chose their answers. As shown in Figure 11, participants given conformal sets chose their answer from the set at an adoption rate very similar to the set’s stated coverage guarantee. Reliance on top-$k$ sets is higher than expected, which is also a result of the larger average set size for top-$k$.
Figure 11: Adoption Rate by Treatment Group.
@@ 6 Conclusion and Outlook
It is easy to assume without further reflection that uncertainty quantification using methods such as conformal prediction will make model predictions more useful to humans. However, in science we must not blindly follow such assumptions without evidence to support them. In this work, we conducted a scientific study to verify the intuitive notion that providing prediction sets from models can improve human decision making. By recruiting 600 paid participants across all tests, and collecting a total of 42,500 individual responses, we found statistically significant evidence that conformal prediction sets can improve human accuracy on classification tasks, both compared to no assistance, and top-$k$ sets. Since we controlled for the level of empirical coverage provided by conformal and top-$k$ sets, we ascribe the improvements in accuracy to two factors: conformal sets have smaller average set sizes, and they quantify model uncertainty via their distribution of set sizes. In contrast to accuracy, we did not find consistent evidence that conformal sets increase decision making speed. A possible explanation is that parsing prediction sets can increase the amount of cognitive processing required to make an informed decision.
@@
Our work informs the design of human-in-the-loop decision pipelines (Wu et al., 2022), which we point out are a natural solution to the problem that set-valued predictions may not be actionable. Including a human into decision pipelines can mitigate some of the concerns around the trustworthiness of machine learning models. For example, more advanced machine learning methods are often less explainable (Linardatos et al., 2021), whereas humans could articulate their thought process even when aided by prediction sets. Models are susceptible to perpetuating the biases of their training data leading to fairness concerns (Mehrabi et al., 2021), whereas humans could consciously override a prediction set when they see unfair outcomes. Neural networks are infamously susceptible to adversarial examples (Biggio et al., 2013; Szegedy et al., 2013; Goodfellow et al., 2014), which, by definition, humans are not fooled by, making a human-in-the-loop perhaps the most robust defense to adversarial attacks.
@@
Finally, while we only covered classification tasks in this study, it would be interesting to examine the compatibility of human decision making with conformal prediction for regression or time series problems. We also only considered tasks that could be completed by the general population. Our society still relies on experts, such as medical doctors, for many critical decisions and is not prepared to accept the risks of automating them. However, there are potentially great benefits of augmenting the skills of experts by providing information from a model in the form of conformal prediction sets, leaving ultimate control in the hands of the experts (Tizhoosh & Pantanowitz, 2018; Kompa et al., 2021).
7 Impact Statement
@@
Our work promotes the integration of humans into decision making pipelines as a means to mitigate negative impacts of untrustworthy machine learning. As such, we do not foresee negative societal impacts of our study. We provide a complete of research ethics for our human subject experiments in Appendix B.
@@ A.1 Dataset Pre-processing
High level statistics about the calibration and test datasets are given in Table 4. Below, we give complete details on the pre-processing applied to each dataset.
@@
Since ObjectNet is comprised of only a test set (50,000 images), we split it to obtain images for calibration. First, we selected the $M=20$ most common classes from all 313 to ensure humans could easily parse all the listed classes. The resulting classes were: [‘Backpack’, ‘Banana’, ‘Bandage’, ‘Battery’, ‘Belt’, ‘Blanket’, ‘Book’, ‘Bottle’, ‘Bottle Cap’, ‘Bottle Opener’, ‘Broom’, ‘Bucket’, ‘Candle’, ‘Cellphone’, ‘Cellphone Charger’, ‘Envelope’, ‘Figurine’, ‘Sandal’, ‘Knife’, ‘Trash Bin’]. Of the remaining samples, we then separated them into calibration and test datasets, reserving 2000 samples for calibration with the remaining samples used for testing. During splitting, we additionally performed stratified sampling on both the calibration and test sets separately to ensure that classes were balanced. As both calibration and test sets were treated the same way, they can be considered i.i.d. for conformal prediction. Finally, each image was resized via bicubic and center-cropping transformations to $224\times 224$ pixels following the image processing pipeline of CLIP (Radford et al., 2021) so that both the model and humans were presented with consistent images.
@@
We used the original validation dataset for conformal calibration and the original test set as our test set. For pre-processing, we kept only the sentences with a single label and removed any sentences containing emojis, as emojis were incompatible with PsychoPy. Then, to make the task more manageable for human evaluators, we picked the $M=10$ most common sentiment classes based on the calibration set [‘Admiration’, ‘Gratitude’, ‘Approval’, ‘Disapproval’, ‘Amusement’, ‘Annoyance’, ‘Curiosity’, ‘Love’, ‘Optimism’, ‘Neutral’] out of the 28 overall classes. Finally we performed stratified sampling on both calibration and test sets so that each selected class had the same amount of samples in both datasets. As both calibration and test sets were treated the same way, they can be considered i.i.d. for conformal prediction. The resulting datasets contained 1180 calibration samples and 1030 test samples.
@@
We used the validation and test sets that are defined in the supervised variant of Few-NERD as our calibration and test sets, respectively. We first filtered out text examples that contained non-ASCII characters because of limitations when displaying through PsychoPy. We also filtered out examples where the longest named entity spans more than 8 words, where the entire example is longer than 60 words, and where the named entity is presented without any context. Then we selected the $M=20$ most common classes which were [‘person-actor’, ‘person-artist/author’, ‘person-athlete’, ‘other-award’, ‘location-bodiesofwater’, ‘other-biologything’, ‘organization-company’, ‘event-attack/battle/war/militaryconflict’, ‘organization-education’, ‘location-GPE’, ‘organization-government/governmentagency’, ‘organization-media/newspaper’, ‘art-music’, ‘organization-politicalparty’, ‘person-politician’, ‘event-sportsevent’, ‘organization-sportsleague’, ‘organization-sportsteam’, ‘location-road/railway/highway/transit’, ‘art-writtenart’]. For presentation to humans we relabeled these classes as [‘Actor/Actress’, ‘Artist/Author’, ‘Athlete’, ‘Award’, ‘Body of Water’, ‘Biology’, ‘Company’, ‘Military Conflict’, ‘Education’, ‘Geopolitics’, ‘Government’, ‘Media Organization’, ‘Music’, ‘Political Party’, ‘Politician/Leader’, ‘Sports Event’, ‘Sports League’, ‘Sports Team’, ‘Transportation Route’, ‘Written Art’]. We performed stratified sampling separately on the validation and test datasets to balance their classes such that a single named entity was selected and highlighted for examples that contained multiple. Hence this became a balanced multi-class classification task. Finally, we reduced the size of the test set because only 2,500 individual trials were planned for each treatment ($N=50$ participants, $m=50$ trials each). Throughout the process the calibration and test sets were treated the same way, so we can consider them i.i.d. for conformal prediction. The final datasets comprised 5000 calibration samples and 2000 test samples.
In our pre-registration at osf.io/fkdhv, we stated that one of the three datasets used would be Broad Twitter Corpus (Derczynski et al., 2016), an NER task gathered from Twitter data. However, upon preparing this dataset for experiments, we determined it to be unsuitable because it contained many non-English example sentences, despite claims in (Derczynski et al., 2016) that it used English language tweets. We replaced Broad Twitter Corpus with another NER dataset, Few-NERD, which contained only English sentences from Wikipedia.

### v3

@@ 4.2 Experiment Design
Our experimental design was informed by research on the collection of high quality data from crowd sourcing platforms. Mitra et al. (2015) show that the most reliable way to ensure high quality data is by training participants on the task at hand, while providing financial incentives can also be beneficial. Hence, in each experiment participants were asked for consent, given instructions on the task, and trained with 20 examples after which the testing phase began. Participants were paid a fixed amount with a financial incentive proportional to their accuracy. On average, participants were paid 7.80 GBP/hr, and in total we spent 1500 GBP on participant compensation, excluding fees.
An example trial screen is shown in Figure 2, while complete descriptions of the tasks are given in Appendix B. On each trial the participant was shown a stimulus $x$ and all $M$ class labels, and was forced to classify $x$ as one of the labels. For the top-$k$ and conformal treatments, they were also shown a prediction set, along with the (empirical) coverage guarantee. The correct answer was displayed once the participant confirmed their decision. There was no time limit to enter responses, although for ObjectNet the stimulus was only shown for 0.22s, within the limits of human perception (Fraisse, 1984), to increase the difficulty of the task. The set of $m$ stimuli shown to each participant was drawn from $\mathcal{D}_{\mathrm{test}}$, making each test randomized.
Figure 2: Main trial screen shown to participants for ObjectNet with conformal set treatment. The correct answer is given only after the participant responds.
@@
Figure 4: Human performance (speed) across three tasks and three treatments. Data is shown as mean response time (s), while error bars show unbiased standard errors ($N=50$).
@@ 5.1 Human Performance Measurement
Table 2: Accuracy – $p$-values and Effect Sizes
ObjectNet GoEmotions Few-NERD
Comparison $p$ $d$ $p$ $d$ $p$ $d$
Top-$k$ $>$ Control $0.1$ 0.3 $5\mathrm{e-}{5}$ 0.8 $0.003$ 0.6
Conf. $>$ Control $5\mathrm{e-}{5}$ 0.8 $5\mathrm{e-}{9}$ 1.0 $3\mathrm{e-}{8}$ 1.0
Conf. $>$ Top-$k$ $0.01$ 0.5 $0.02$ 0.4 $8\mathrm{e-}{4}$ 0.7
Table 3: Response Time – $p$-values and Effect Sizes
ObjectNet GoEmotions Few-NERD
Comparison $p$ $d$ $p$ $d$ $p$ $d$
Top-$k$ vs. Control $0.6$ 0.1 $0.1$ 0.3 $0.3$ 0.2
Conf. vs. Control $0.07$ 0.4 $0.3$ 0.2 $6\mathrm{e-}{5}$ 0.9
Conf. vs. Top-$k$ $0.2$ 0.2 $0.5$ 0.1 $4\mathrm{e-}{5}$ 0.9
For accuracy, we see that top-$k$ sets mostly do improve human performance compared to the control, while conformal sets lead to better performance than top-$k$ sets. To confirm these visual trends, we conducted pre-planned comparisons between pairs of treatments using $t$-tests as described in Section 3. In all cases except one we reject the null hypothesis that the mean accuracy is the same between treatments, using the significance threshold $p<0.05$ (Table 2, non-significant results in red). Notably, for all three independent tasks we find statistically significant evidence that conformal prediction sets are more useful to humans than top-$k$ sets, with medium effect sizes ($0.4\leq d\leq 0.7$) (Cohen, 2013). Because both types of prediction sets have the same (empirical) coverage, there are only two differences between the methods to which we can ascribe the improvement: conformal sets are smaller on average (Table 1), and conformal sets quantify uncertainty.
We also pre-planned comparisons between the treatments on the average response time. However, in viewing Figure 4 we do not see a consistent trend between all tasks. Conformal sets have the lowest average response time on two tasks, while the control treatment led to the fastest completions on GoEmotions. In most cases we do not reject the null hypothesis that the treatment has no effect on the mean response time (Table 3). The additional information provided as a prediction set must be processed by the user and incorporated into decision making, which can outweigh the speed advantages of receiving a curated shortlist.
@@ 5.2 Ablations
Figure 7: Comparison of ObjectNet results between two models. The ViT-L/14 model achieved 83.3% top-1 accuracy, which was significantly better than ViT-B/32 at 61.1% on ObjectNet test data $\mathcal{D}_{\mathrm{test}}$.
@@
Examining the set size distribution between the two settings helps reveal the source of performance differences. As shown in Figure 5, human performance decreases on examples where prediction sets were larger, but importantly human performance conditioned on set size was comparable between the settings. By comparison, Figure 6 shows that the large setting generated fewer singleton sets where human accuracy is highest, resulting in the observed overall decrease in human accuracy. Optimizing for set sizes where human performance is highest is one way to improve the usefulness of conformal sets for humans.
@@
By swapping the ViT-L/14 model used in Section 5.1 for a weaker ViT-B/32, we can study the effect of model performance on the usefulness of both top-$k$ and conformal prediction sets. 7(a) shows conformal sets greatly enhanced human performance when using a superior underlying model, whereas for top-$k$ the improved model did not translate to as much human improvement. This is despite the empirical coverage (7(b)) increasing by a similar amount between the treatments (by design of our calibration). Furthermore, the rate at which humans chose an answer from the prediction set, which we call the adoption rate (7(c)), was consistently close to the empirical coverage (coverage was communicated to participants during the test). The outsized increase in accuracy for conformal when the better model was used is a function of the higher quality of sets that it produces; although coverage and adoption rates are the same, humans must still choose the correct answer out of the set when they adopt from it, which is much easier to do for singleton sets. The stronger model produces a distribution of conformal set sizes more skewed towards singleton sets (7(d)), whereas top-$k$ sets always use $k=3$ even with a more confident and correct model.
@@ 5.3 Insights
In Figure 3 we concluded that the combination of human and model (conformal treatment) outperforms humans alone (control) in terms of accuracy. Another relevant question is whether human-model teams outperform the model alone. From Table 1 we see that all treatments outperformed the top-1 accuracy of the model on ObjectNet, but the reverse is true on GoEmotions and Few-NERD. In Figure 10 we examine this comparison in more detail using per-class accuracies for ObjectNet. We see high variance in the model’s per-class accuracy, possibly due to the zero-shot nature of predictions we implemented, whereas humans were more consistent (standard deviation $13.4$ vs. $7.8$). Notably, low model accuracy (Bandage, Battery, Blanket, Bucket, Cellphone) tended to drag down the human-model team compared to humans alone. This is a point of caution for implementing human-in-the-loop systems, as poor model performance or biases against certain groups may not be completely overcome by humans. However, when the model outperformed the humans, the human-model team also tended to benefit, sometimes surpassing both individual partners (Backpack, Book, Bottle, Opener, Candle, Sandal). Hence there is a type of ensembling effect (Dietterich, 2000) where one partner’s weaknesses can be corrected by the other.
@@
Participants in the top-$k$ and conformal treatment groups were informed that the prediction sets they were shown contained the true answer with probability $1-\hat{\alpha}$, but were not constrained in how they chose their answers. As shown in Table 4, participants given conformal sets chose their answer from the set at an adoption rate very similar to the set’s stated coverage guarantee, whereas reliance on top-$k$ sets was higher than expected. This may suggest that the distributional information of variable-sized conformal sets better communicates when sets can be trusted.
Table 4: Adoption Rate (%) by Treatment Group
Task $1-\hat{\alpha}$ Top-$k$ Conformal
ObjectNet 94 97 95
GoEmotions 92 94 92
Few-NERD 98 99 99
@@ 6 Conclusion and Outlook
It is easy to assume without further reflection that uncertainty quantification through conformal prediction will make model predictions more useful to humans. However, in science we must not blindly follow such assumptions without evidence to support them. In this work, we conducted a scientific study to verify the intuitive notion that providing prediction sets from models can improve human decision making. By recruiting 600 paid participants across all tests, and collecting a total of 42,500 individual responses, we found statistically significant evidence that conformal prediction sets can improve human accuracy on classification tasks, both compared to no assistance, and to top-$k$ sets. Since we controlled for the level of empirical coverage provided by conformal and top-$k$ sets, we ascribe the improvements in accuracy to two factors: conformal sets have smaller average set sizes, and they quantify model uncertainty via their distribution of set sizes. In contrast to accuracy, we did not find consistent evidence that conformal sets increase decision making speed. A possible explanation is that parsing prediction sets can increase the amount of cognitive processing required to make an informed decision.
@@
Our work informs the design of human-in-the-loop decision pipelines (Wu et al., 2022), which we point out are a natural solution to the problem that set-valued predictions may not be actionable. Including a human into decision pipelines can mitigate some of the concerns around the trustworthiness of machine learning models. For example, more advanced machine learning methods are often less explainable (Linardatos et al., 2021), whereas humans could articulate their thought process even when aided by prediction sets. Neural networks are infamously susceptible to adversarial examples (Biggio et al., 2013; Szegedy et al., 2013; Goodfellow et al., 2014), which, by definition, do not fool humans, making a human-in-the-loop perhaps the most robust defense to adversarial attacks. Still, there are limitations to human-in-the-loop systems. On two out of three tasks, top-1 model accuracy was higher than what humans achieved (even with model assistance). We also found that when a model performs particularly poorly on groups within the data, prediction sets can drag down human performance on those groups. This could manifest as a transfer of biases from models to humans, and reinforces that the fairness of models still needs to be considered even with a human-in-the-loop (Mehrabi et al., 2021).
@@
Finally, while we only covered classification tasks in this study, it would be interesting to examine the compatibility of humans with conformal prediction for regression and time series problems. We also only considered tasks that could be completed by the general population. Our society still relies on experts, such as medical doctors, for many critical decisions and is not prepared to accept the risks of automating them. However, there are potentially great benefits to augmenting the skills of experts by providing information from a model in the form of conformal prediction sets, leaving ultimate control in the hands of the experts (Tizhoosh & Pantanowitz, 2018; Kompa et al., 2021).
Impact Statement
@@
Our work promotes the integration of humans into decision making pipelines as a means to mitigate negative impacts of untrustworthy machine learning. As such, we do not foresee negative societal impacts of our study. We provide a complete description of research ethics for our human subject experiments in Appendix B.
Acknowledgements
We would like to thank Mouloud Belbahri and Brendan Ross for comments on a draft of this manuscript.
@@ A.1 Dataset Pre-processing
High level statistics about the calibration and test datasets are given in Table 5. Below, we give complete details on the pre-processing applied to each dataset. All pre-processing including calibration and generation of conformal sets was carried out with an Intel Xeon Silver 4114 CPU and TITAN V GPU and takes under 3 hours. We provide the code to generate all datasets at github.com/layer6ai-labs/hitl-conformal-prediction.
@@
Since ObjectNet is comprised of only a test set (50,000 images), we split it to obtain images for calibration. First, we selected the $M=20$ most common classes from all 313 to ensure humans could easily parse all the listed classes. The resulting classes were: [‘Backpack’, ‘Banana’, ‘Bandage’, ‘Battery’, ‘Belt’, ‘Blanket’, ‘Book’, ‘Bottle’, ‘Bottle Cap’, ‘Bottle Opener’, ‘Broom’, ‘Bucket’, ‘Candle’, ‘Cellphone’, ‘Cellphone Charger’, ‘Envelope’, ‘Figurine’, ‘Sandal’, ‘Knife’, ‘Trash Bin’]. We then separated the selected samples into calibration and test datasets, reserving 2000 samples for calibration with the remaining samples used for testing. During splitting, we additionally performed stratified sampling on both the calibration and test sets separately to ensure that classes were balanced. As both calibration and test sets were treated the same way, they can be considered i.i.d. for conformal prediction. Finally, each image was resized via bicubic and center-cropping transformations to $224\times 224$ pixels following the image processing pipeline of CLIP (Radford et al., 2021) so that both the model and humans were presented with consistent images. ObjectNet is released under a license that permits free use for research purposes.
@@
We used the original validation dataset for conformal calibration and the original test set as our test set. For pre-processing, we kept only the sentences with a single label and removed any sentences containing emojis, as emojis were incompatible with PsychoPy. Then, to make the task more manageable for human evaluators, we picked the $M=10$ most common sentiment classes based on the calibration set [‘Admiration’, ‘Gratitude’, ‘Approval’, ‘Disapproval’, ‘Amusement’, ‘Annoyance’, ‘Curiosity’, ‘Love’, ‘Optimism’, ‘Neutral’] out of the 28 overall classes. Finally we performed stratified sampling on both calibration and test sets so that the selected classes were balanced. As both calibration and test sets were treated the same way, they can be considered i.i.d. for conformal prediction. The resulting datasets contained 1180 calibration samples and 1030 test samples. GoEmotions is released under a license that permits free use for research purposes.
@@
We used the validation and test sets that are defined in the supervised variant of Few-NERD as our calibration and test sets, respectively. We first filtered out text examples that contained non-ASCII characters because of limitations when displaying through PsychoPy. We also filtered out examples where the longest named entity spans more than 8 words, where the entire example is longer than 60 words, and where the named entity is presented without any context. Then we selected the $M=20$ most common classes which were [‘person-actor’, ‘person-artist/author’, ‘person-athlete’, ‘other-award’, ‘location-bodiesofwater’, ‘other-biologything’, ‘organization-company’, ‘event-attack/battle/war/militaryconflict’, ‘organization-education’, ‘location-GPE’, ‘organization-government/governmentagency’, ‘organization-media/newspaper’, ‘art-music’, ‘organization-politicalparty’, ‘person-politician’, ‘event-sportsevent’, ‘organization-sportsleague’, ‘organization-sportsteam’, ‘location-road/railway/highway/transit’, ‘art-writtenart’]. For presentation to humans we relabeled these classes as [‘Actor/Actress’, ‘Artist/Author’, ‘Athlete’, ‘Award’, ‘Body of Water’, ‘Biology’, ‘Company’, ‘Military Conflict’, ‘Education’, ‘Geopolitics’, ‘Government’, ‘Media Organization’, ‘Music’, ‘Political Party’, ‘Politician/Leader’, ‘Sports Event’, ‘Sports League’, ‘Sports Team’, ‘Transportation Route’, ‘Written Art’]. We performed stratified sampling separately on the validation and test datasets to balance their classes such that a single named entity was selected and highlighted for examples that contained multiple. Hence this became a balanced multi-class classification task. Finally, we reduced the size of the test set because only 2,500 individual trials were planned for each treatment ($N=50$ participants, $m=50$ trials each). Throughout the process the calibration and test sets were treated the same way, so we can consider them i.i.d. for conformal prediction. The final datasets comprised 5000 calibration samples and 2000 test samples. Few-NERD is released under a license that permits free use for research purposes.
In our pre-registration, we stated that one of the three datasets used would be Broad Twitter Corpus (Derczynski et al., 2016), an NER task gathered from Twitter data. However, upon preparing this dataset for experiments, we determined it to be unsuitable because it contained many non-English example sentences, despite claims in (Derczynski et al., 2016) that it used English language tweets. We replaced Broad Twitter Corpus with another NER dataset, Few-NERD, which contained only English sentences from Wikipedia.